a certain disease haa a mortality rate of 75%. Two patients suffering from the disease are selected at random. What is the probability that at least one of them will recover?
1
Expert's answer
2012-10-22T12:17:18-0400
probability that at least one of them will recover = 1 - probability that both
are going to die = 1 - 3/4 * 3/4 = 1-9/16 = 0.4375